Tom's present age is twice Sue's present age. Two years ago, Tom was three times as old as Sue was then. Find their present ages.
Answer by mananth(16946)

let sue's age be x
tom's age will be 2x
2 years ago sue was x-2 years
tom was 2x-2 years
At that time tom was thrice as old as sue
3(x-2) = 2x-2
3x-6 = 2x-2
x= 6-2
x=4 which is sue's age
tom is 2x = 8 years